how do I solve:
x2+5=30 and
x2+x+30
Please tell me the answers A.S.A.P

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

how do I solve:
x2+5=30
Subtract 5, take the square root.
=================
and
x2+x+30
No equal sign, not an equation, nothing to solve.

Answer by stanbon(75887)

how do I solve:
x2+5=30
x^2 = 25
x = +5 or x = -5
---------------------
x2+x+30 =
x = [-1 +- sqrt(1-4*1*30)]/2
----
x = [-1 +- sqrt(-119)]/2
----
x = [-1 +- *sqrt(119)i]/2
===============================
Cheers,
Stan H.
============